In the market for Fante Kenley, the supply and demand functions respectively are Qs = 0.25P+10 and  Qd = -0.5P+100

When there is excess demand, price adjusts according to the equation dp/dt = 0.5 (Qd - Qs)

a) Find the long run equilibrium price, P* (that is, the price at which there is no excess demand or supply). 

b) Formulate and solve he first order differential equation giving P as a function of time, t. Is this market dynamically stable or unstable? 

c) If the initial price is P = 50, how close will the price be to its long run equilibrium value, when t = 10? 



1
Expert's answer
2021-08-02T02:56:02-0400

a)

"Qs = 0.25P + 10\\

Qd = -0.5p + 100"

If excess demand , price adjustment "\frac{∂p}{∂t} = 0.5(Q^d − Q^s)"

For Long-run equilibrium put

"QD = Qd\\

0.25P + 10 = -0.5p + 100\\

0.25P + 0.5p = 100 - 10

0.30P = 90\\

P = \frac{90}{0.30}\\

P = 300"

Now put P = 300 in any equation to calculate the Equilibrium Quantity

Qs = 0.25P + 10

Qs = 0.25*300 + 10

Qs = 75 + 10

Qs = 80

"Qs = 0.25P + 10\\

Qs = 0.25\times300 + 10\\

Qs = 75 + 10\\

Qs = 80"


b)

The Given Differentiation is "\frac{∂p}{∂t} = 0.5(Q^d − Q^s)"

"\frac{∂p}{∂t} = 0.5(-.5p+100-0.25p-10)"

which can be arranged as

"\frac{∂p}{∂t} = 0.5(-.5p+100-0.25p-10)"

"\frac{∂p}{∂t} = -0.15p+45" where b=-0.15 and c=45

"p=Ae^{bt}-\frac{c}{b}" where "A=P(0)+\frac{c}{b}"

since b=-0.15 and c=45, we have "\frac{c}{b}=-p^*" (i.e equilibrium price)

"P = (P(0) + P*)e^{bt}+p^*"

"P = (P(0) + 300)e^{0.15t} + 300"

 Accordingly as b is < 0 this implies it will increase monotonically

 through time, therefore market is Dynamically unstable


c)

"When\space P(0) = 50, A\\

P= (P(0) + 300)_e^{0.15t}− 300\\P = (50 + 300)_e^{0.15\times10} − 300\\P = (350)\times4.4816 − 300\\P = 820.422"


Check the explanation abaove